# Multi-Stage Manipulation with Demonstration-Augmented Reward, Policy, and World Model Learning

### v1

@@ 1 Introduction
Prior work shows that Learning from Demonstrations (LfD) helps mitigate exploration issues in sparse reward settings. Algorithms such as CoDER (Zhan et al., 2022) and MoDem (Hansen et al., 2023; Lancaster et al., 2024) leverage demonstrations to populate the replay buffer of an off-policy RL algorithm (Sutton & Barto, 2018), but they often scale poorly with task complexity as they need demonstrations that sufficiently cover the behavior space. Inverse RL methods (Trott et al., 2019; Wu et al., 2021; Memarian et al., 2021; Escontrela et al., 2022) train RL on a reward function that is learned from demonstrations, but inverse RL alone often struggles as the learned reward function can have poor reward predictions on unseen states. Lastly, these methods typically require a vast amount of samples to learn a reward function (Kumar et al., 2022; Mu et al., 2024) before any policy learning can begin.
@@
    We analyze the relative importance of each component of our framework, and are committed to open-sourcing all code and demonstrations used in this work.
@@ 3.1 Model-based RL with online reward learning
Sparse rewards are a major challenge in RL, particularly for long-horizon tasks comprising multiple stages. To overcome this, we learn to densify sparse rewards with a small number of demonstrations. For this, we introduce a series of discriminators $\{\delta_{k}\}_{k=0}^{N}$, each corresponding to a task stage $k\in\{0\dots N\}$. Each discriminator is trained to predict the likelihood of progressing to the next stage based on the latent state representation $\mathbf{z_{t}}$ produced by the world model.
@@ 3.2 Training scheme
[…]
4 Experiments
We consider 16 challenging visual multi-stage manipulation tasks with a long-horizon for our experimental evaluation. This includes 5 manipulation tasks from ManiSkill3 (Tao et al., 2024b), 5 manipulation tasks from Meta-World (Yu et al., 2021) and 4 manipulation tasks from Robosuite (Zhu et al., 2022). Additionally, we include 2 humanoid manipulation tasks from ManiSkill3 with a high-dimensional action space, which we refer to as ManiSkill Humanoids in our evaluations. We place a strong emphasis on long-horizon precise manipulation, which is why we select the most challenging tasks from each domain. We relate the difficulty to the required precision of a task, its horizon, and the level of randomization in the scene (see Appendix D.3 for further details on difficulty categorization). For each task, the agent is given a constrained budget of demonstrations and interaction steps (see Table 1). To allow most baselines to solve the task, we set the interaction budget and demonstrations to a different amount for each task. For a complete list of details on our experimental setup, please refer to Table 1 and Appendix D. We have open-sourced all code and environments111 https://github.com/adrialopezescoriza/demo3. Through our evaluations, we aim to answer the following questions:
@@ 4.1 Baselines
MoDem (Hansen et al., 2023) is an MBRL algorithm designed to enhance data-efficiency in visual control tasks with sparse rewards. Similarly to our method, MoDem employs a three-phase framework: policy pretraining, seeding, and interactive learning with oversampling of demonstration data. The authors show state-of-the-art performance on Meta-World and Adroit domains from visual inputs.
@@ 4.3 Analysis
Figure 7 shows the effect of removing dense reward learning, policy pretraining, and demonstration oversampling in the 5 manipulation tasks from ManiSKill3. Interestingly, a considerable jump in performance is brought by pretraining and oversampling from the demonstration dataset with the TD-MPC2 backbone (no learned reward). The effect of reward learning becomes evident in long-horizon tasks where advancing stages can be very challenging without rewards (e.g., Peg Insertion, Pick Place), especially when reducing the number of demonstrations (see Appendix A.2).
@@
Figure 9 illustrates the data-efficiency of our method across the same tasks under varying reward granularities: 1 stage, 2 stages, 3 stages, and dense rewards. Please refer to D for details about the stage definitions. As expected, dense rewards provide the optimal learning signal, resulting in the fastest task completion. Remarkably, our method achieves near-equal performance with only 2 stages, demonstrating that stage indicators combined with a minimal set of demonstrations are sufficient to guide the learning process effectively. This result shows that over-engineered dense reward functions can easily be avoided by only providing a small amount of demonstrations. By relying solely on stage indicators and demonstrations, our method simplifies the reward design process while maintaining high performance in complex long-horizon tasks.
@@ 6 Conclusions and Future Directions
In this work, we address the challenge of learning long-horizon manipulation skills with sparse rewards using only proprioceptive and visual feedback. We propose name, a demonstration-augmented MBRL algorithm that simultaneously learns a reward function, policy, and world model for multi-stage manipulation tasks. Our experiments (see Section 4) show that name outperforms the state-of-the-art by 40% and converges 4$\times$ faster on the most difficult tasks.
Regarding future work, we identify two key directions. First, while our results indicate that a very small number of demonstrations is sufficient for accelerated learning, we do not investigate the impact of different demonstration sources, e.g., human teleoperation. Although prior work suggests that such variations have a minor effect on RL performance (Hansen et al., 2023), making demonstration-augmented RL methods like name more robust to diverse data sources remains a valuable research direction.
Second, the integration of demonstrations within the MBRL framework warrants further exploration. Our results show that a simple 50% sampling schedule between demonstrations and the replay buffer already provides substantial benefits. This raises the question of whether more advanced sampling strategies could further improve data efficiency. Inspired by prioritized replay buffers (Schaul et al., 2016), optimizing demonstration sampling to maximize learning efficiency presents a promising avenue for future research.
@@ Appendix A Additional Results
A.1 Single-task Experimental Results
@@ D.2 Stage Definitions
Stack Cube Red cube grabbed. Red cube is above green cube. Red cube is placed on green cube and released.
Peg Insertion Peg grabbed. Peg is aligned with hole. Peg inserted in hole.
Pick Place Cube grabbed. Cube is close to goal location. Cube located at goal position and static.
Poke Cube Peg grabbed. Peg is in contact with Cube. Cube is at goal location.
Lift Peg Upright Peg grabbed. Peg is close to upright position. Peg is in the upright position on the desk.
@@ E.2 Hyperparameters
Initial pretrained action ratio $\alpha_{0}$ $5\times 10^{-5}$
@@ E.3 Computational Resources
All our experiments run in a single NVIDIA GeForce RTX 3090 GPU and 32GB of RAM to store collected samples.

### v2

@@ 1 Introduction
Prior work shows that Learning from Demonstrations (LfD) can help mitigate exploration issues in sparse reward settings. Algorithms such as CoDER (Zhan et al., 2022) and MoDem (Hansen et al., 2023; Lancaster et al., 2024) leverage demonstrations to populate the replay buffer of an off-policy RL algorithm (Sutton & Barto, 2018), but they often scale poorly with task complexity as they need demonstrations that sufficiently cover the behavior space. Inverse RL methods (Trott et al., 2019; Wu et al., 2021; Memarian et al., 2021; Escontrela et al., 2022) train RL on a reward function that is learned from demonstrations, but inverse RL alone often struggles as the learned reward function can have poor predictions on unseen states. Lastly, these methods typically require a vast amount of samples to learn a reward function (Kumar et al., 2022; Mu et al., 2024) before any policy learning can begin.
@@
    We analyze the relative importance of each component of our framework, and are open-sourcing all code and demonstrations used in this work.
@@ 3.1 Model-based RL with online reward learning
Sparse rewards are a major challenge in RL, particularly for long-horizon tasks comprising multiple stages. To overcome this, we learn to densify sparse rewards with a small number of demonstrations. For this, we introduce a series of discriminators $\{\delta_{k}\}_{k=0}^{N}$, each corresponding to a task stage $k\in\{0\dots N\}$. The objective of each discriminator is to predict the likelihood of progressing to the next stage based on the latent state representation $\mathbf{z_{t}}$ produced by the back-bone world model.
@@ 4 Experiments
We consider 16 challenging visual multi-stage manipulation tasks with a long-horizon for our experimental evaluation. This includes 5 manipulation tasks from ManiSkill3 (Tao et al., 2024b), 5 manipulation tasks from Meta-World (Yu et al., 2021) and 4 manipulation tasks from Robosuite (Zhu et al., 2022). Additionally, we include 2 humanoid manipulation tasks from ManiSkill3 with a high-dimensional action space, which we refer to as ManiSkill Humanoids in our evaluations. We place a strong emphasis on long-horizon precise manipulation, which is why we select the most challenging tasks from each domain. We relate the difficulty to the required precision of a task, its horizon, and the level of randomization in the scene (see Appendix D.3 for further details on difficulty categorization). For each task, the agent is given a constrained budget of demonstrations and interaction steps (see Table 1). To allow most baselines to solve the task, we set the interaction budget and demonstrations to a different amount for each task. For a complete list of details on our experimental setup, please refer to Table 1 and Appendix D. Through our evaluations, we aim to answer the following questions:
@@ 4.1 Baselines
MoDem (Hansen et al., 2023) is a MBRL algorithm designed to enhance data-efficiency in visual control tasks with sparse rewards. Similarly to our method, MoDem employs a three-phase framework: policy pretraining, seeding, and interactive learning with oversampling of demonstration data. The authors show state-of-the-art performance on Meta-World and Adroit domains from visual inputs and sparse rewards.
@@ 4.3 Analysis
. Figure 7 shows the effect of removing dense reward learning, policy pretraining, and demonstration oversampling in the 5 manipulation tasks from ManiSkill3. Interestingly, a considerable jump in performance is brought by pretraining and oversampling from the demonstration dataset with the TD-MPC2 backbone (no learned reward). The effect of reward learning becomes evident in long-horizon tasks where advancing stages can be very challenging without rewards (e.g., Peg Insertion, Pick Place), especially when reducing the number of demonstrations (see Appendix A.2).
@@
Figure 9 illustrates the data-efficiency of our method across the same tasks under varying reward granularities: 1 stage, 2 stages, 3 stages, and dense rewards. Please refer to D for details about the stage definitions. As expected, dense rewards provide the optimal learning signal, resulting in the fastest task completion. Remarkably, our method achieves near-equal performance with only 2 stages, demonstrating that stage indicators combined with a minimal set of demonstrations are sufficient to guide the learning process effectively. This result shows that over-engineered dense reward functions can easily be avoided by only providing a small number of demonstrations. By relying solely on stage indicators and demonstrations, our method simplifies the reward design process while maintaining high performance in complex long-horizon tasks.
@@ 6 Conclusions and Future Directions
[…]
Firstly, although our experiments show that a small number of demonstrations is sufficient for accelerated learning, we do not explore the effect of using different sources, e.g., data collected via human teleoperation. While prior work suggests such sources have limited impact on RL performance (Hansen et al., 2023), improving robustness to diverse demonstration types remains relevant for deployment.
Another limitation is that our experiments are limited to simulation. We aim to deploy name on real robot hardware to evaluate its generalization under domain shift and whether the same constrained regime of demonstrations and interaction samples can yield reliable real-world policies. Notably, Hansen et al. (2023) shows that a similar pipeline transfers well to physical systems. Since name improves on these components in simulation, we are optimistic its benefits will carry over.
[…]
Acknowledgements
Nicklas Hansen is supported by NVIDIA Graduate Fellowship. Stone Tao is supported in part by the NSF Graduate Research Fellowship Program grant under grant No. DGE2038238.
Impact Statement
This paper presents work whose goal is to advance the field of robot learning. While there are many potential societal consequences in developing embodied intelligence for the real world, we do not feel our work presents any particular implications that should be highlighted here.
@@ Appendix A Additional Results
A.1 Single Task Experimental Results
@@ A.2 Additional Ablations
A.3 Learned Reward
Figure 15: Learned reward animation. Visualization of the learned dense reward across three representative rollouts—successful, failed, and semi-successful. The reward curves closely track task progress.
@@ D.2 Stage Definitions
Stack Cube Cube A grabbed. Cube A above Cube B. Cube A stacked on top of Cube B.
Peg Insertion Peg grabbed. Peg aligned with hole. Peg inserted in hole.
Pick Place Cube grabbed. Cube close to goal. Cube at goal.
Poke Cube Peg grabbed. Peg touching Cube. Cube at goal.
Lift Peg Upright Peg grabbed. Peg upright. Peg upright on desk.
@@ E.2 Hyperparameters
$\alpha_{0}$ $5\times 10^{-5}$
@@ E.3 Computational Resources
All our experiments run on a single NVIDIA GeForce RTX 3090 GPU and 32GB of RAM to store collected samples.